### v1

@@ 5.3 RQ2: Performance of prompt engineering
5.4 RQ3: Selection of classification model
The detection capability of AppPoet relies not solely on the quality of prompt engineering, selecting appropriate classification model is of equal importance. To determine the suitable model for classification detection, model selection experiments are conducted. This experiment broadly considers various representative models, which is consistent with the experimental setup outlined in Section 5.2. Specifically, five typical classification models are selected: CNN (Simonyan and Zisserman (2014)), TextCNN (Chen (2015)), RNN (Elman (1990)), LSTM (Hochreiter and Schmidhuber (1997)), and MLP (Schmidhuber (2015)). The structure and parameters of these models are tuned and then their classification capabilities are evaluated and compared. The experimental outcomes are presented in Table 9. It is worth noting that the selected models require different input formats, thus the inputs need to be adjusted to match the format required for each specific model selected. For example, the CNN model utilizes a representation matrix, while the MLP model requires the matrix to be transformed into a sequential representation vector before input.
Table 9: Comparison of different classification model.
Method ACC(%) Precision(%) Recall(%) F1(%)
CNN 95.79 96.04 95.71 95.87
TextCNN 96.03 96.05 96.17 96.11
RNN 95.94 96.92 95.08 95.99
LSTM 96.07 96.68 95.59 96.13
MLP 97.15 97.03 97.39 97.21
According to the results, the MLP model is identified as the most suitable for the detection task, outperforming all other models across all metrics. A plausible explanation for this is the use of the advanced embedding model text-embedding-ada-002 to obtain representation vectors, which already possess strong representational capabilities. Classification models such as CNN and RNN may somewhat degrade the semantic information that has already been captured, whereas MLP’s approach to dimensionality reduction in classification maximizes information retention, leading to excellent classification outcomes. It should be noted that our current considerations in selecting classification models are all well-established classical algorithms. In the future, we will introduce more advanced algorithms for classification.
5.5 RQ4: Performance of diagnostic report
@@
Specifically, a malicious application with the package name "com.applucinante.weddingrings" is randomly selected for analysis, using AppPoet and Drebin Arp et al. (2014). Both methods accurately identify the application as malware. Then, we generate an interpretable report using the two methods respectively, as shown in Appendix B. It should be noted that we manually verify the application after decompiling it. According to the information provided in the report, we locate that the App has a privacy leakage of obtaining sensitive information about the phone and exporting it to outside.
@@ 6 Discussion
In this section, we discuss the current limitations in AppPoet and the selection and use of LLM.
Current limitations in AppPoet’s implementation. In Section 4.1, we completed our work by selecting several features that are more typical of feature engineering methods. In fact, our approach is highly scalable and can continuously integrate more static features to combine into different views. In the future, we will continue to expand the richness of the views to enhance the accuracy and generalization of detection. In Section 4.2, we design rigorous prompt templates to guide the LLM in outputting the specified content according to the requirements and formats. Although we verified the excellent performance of AppPoet in a real application dataset, considering the possibility of LLM’s hallucinations and the rigor of the system design, an effective mechanism for checking and correcting errors needs to be established in the next step. In addition, our method is based exclusively on the static analysis of APK internal information, which may introduce a certain amount of false positives in our detection results and diagnostic reports. Therefore, an LLM-driven interactive dynamic and static combined detection method is our next step in this direction.
@@ 7 Conclusion
With the rapid development of the Android operating system, Android malware detection has emerged as a critical issue within the community. Although many existing feature engineering-based methods yield good results, they predominantly face two challenges: insufficient depth in feature mining and the difficulty of producing readable diagnostic reports. Inspired by the success of LLM in natural language understanding, we introduce AppPoet, a novel prompt engineering-based method for Android malware detection. Specifically, we select permission, API, URL, and use-feature as entry points for observing Android applications, combining them into three independent views. We then direct the LLM to generate function descriptions and view summaries for each view through our proposed multi-view prompt engineering method. Subsequently, this textual information is converted into machine-readable representation vectors, enabling malware identification through a trained DNN classifier. Finally, we utilize the discrimination results and the descriptive and summary texts of each view to direct the LLM in generating a diagnostic report on the application. This report serves as a guide for reviewing and further analyzing the problem. Through the aforementioned method, it becomes possible to further mine behavioral information concealed within the features, to achieve more precise malware detection via the fusion of multi-view information, and to produce a user-friendly diagnostic report.

### v3

@@ 5.3 RQ2: Performance of prompt engineering
Table 12: Average detection time consumption.
Phase of the detection Average time consumption (s)
Feature extraction 2.31
Function description generation 0.57
View summary generation 8.77
Embedding and detection 2.37
Total 14.02
To further investigate the real-world efficiency of AppPoet, we train a classifier on the small-scale dataset used in the memory component experiments. We then introduce 1,000 samples (an equal split between malicious and benign) from the larger dataset as unknown APK inputs. Starting from the feature extraction step, we record the average consumption at each stage of the detection process, as shown in Table 12. It is important to note that the memory generated during the function description generation phase for the training samples is reused in the process of generating related texts for unknown samples. Additionally, since the text generation process for each view is entirely independent, we implement a multi-threading approach in the real detection environment to handle different views in parallel. The results show that AppPoet takes an average of 14.02s to detect a single APK, achieving 93.7% ACC and 93.64% F1 score. While maintaining high detection accuracy, it also demonstrates good efficiency, making it capable of real-time detection and scalability in real-world applications.
5.4 RQ3: Performance of diagnostic report
@@
Specifically, a malicious application with the package name "com.applucinante.weddingrings" is randomly selected for analysis, using AppPoet and Drebin (Arp et al., 2014). Both methods accurately identify the application as malware. Then, we generate an interpretable report using the two methods respectively, as shown in Appendix B. It should be noted that we manually verify the application after decompiling it. According to the information provided in the report, we locate that the App has a privacy leakage of obtaining sensitive information about the phone and exporting it to outside.
@@ 6 Discussion
In this section, we discuss the current limitations in AppPoet as well as the selection and use of LLM.
Current limitations in AppPoet’s implementation. In Section 4.1, we completed our work by selecting several features that are more typical of learning-based methods. In fact, our approach is highly scalable and can continuously integrate more static features to combine into different views. In the future, we will continue to expand the richness of the views to enhance the accuracy and generalization of detection. In Section 4.2, we design rigorous prompt templates to guide the LLM in outputting the specified content according to the requirements and formats. Although we verified the excellent performance of AppPoet in a real application dataset, considering the possibility of LLM’s hallucinations and the rigor of the system design, an effective mechanism for checking and correcting errors needs to be established in the next step. In addition, our method is based exclusively on the static analysis of APK internal information, which may introduce a certain amount of false positives in our detection results and diagnostic reports. Therefore, an LLM-driven interactive dynamic and static combined detection method is our next step in this direction. Considering that malware is continuously evolving, we also plan to regularly update and expand the dataset in the future to enhance AppPoet’s robustness.
@@ 7 Conclusion
With the rapid development of the Android operating system, Android malware detection has emerged as a critical issue within the community. Existing String-based and Image-based methods generally lack the mining of semantic information about feature behavior, which affects the detection accuracy of the methods to some extent. While graph-based methods are able to mine implicit semantics of features by constructing graphs, they often imply high complexity and abstraction. Inspired by the success of LLM in natural language understanding, we introduce AppPoet, a novel prompt engineering-based method for Android malware detection. Specifically, we select permission, API, URL, and uses-feature as entry points for observing Android applications, combining them into three independent views. We then direct the LLM to generate function descriptions and view summaries for each view through our proposed multi-view prompt engineering method. Subsequently, this textual information is converted into machine-readable representation vectors, enabling malware identification through a trained DNN classifier. Finally, we utilize the discrimination results and the descriptive and summary texts of each view to direct the LLM in generating a diagnostic report on the application. This report serves as a guide for reviewing and further analyzing the problem. Through the aforementioned method, it becomes possible to further mine behavioral information concealed within the features, to achieve more precise malware detection via the fusion of multi-view information, and to produce a user-friendly diagnostic report.
Acknowledgements
This work is supported in part by Anhui Province Natural Science Foundation under Grant No.2408085MF167 and No.2108085QF262, National Natural Science Foundation of China under Grant No.62102385. We thank all the anonymous reviewers who generously contributed their time and efforts. Their professional recommendations have greatly enhanced the quality of the manuscript.